### v1

# Intelligence Preschool Education System based on Multimodal Interaction Systems and AI

Intelligence Preschool Education System based on Multimodal Interaction Systems and AI
Long Xu
(July 2024)
Author Long Xu
1 Introduction
Rapid progress in AI technologies has generated considerable interest in their potential to address challenges in every field and education is no exception. Improving learning outcomes and providing relevant education to all have been dominant themes universally, both in the developed and developing world. And they have taken on greater significance in the current era of technology driven personalization.
1.1 Research Motivation
Although the Chinese government has invested heavily in education and has made great progress in all aspects, it still faces huge challenges in the field of preschool education.
Early childhood education is of great significance in fostering children’s holistic development, forming habits, enhancing cognition, and reducing developmental gaps, etc. However, the current early childhood education model has the following two shortcomings.
  • The quality of education varies widely across different regions, leading to educational inequity.
  • The educational process is difficult to digitize, and the outcomes tend to rely more on subjective perception. This makes it challenging to promote effective educational methods and prevents the improvement of inadequate ones.
Inspired by the development of artificial intelligence and the internet, we propose a digital early childhood education system based on these technologies to address the first issue. For the second issue, we will focus on researching the key factors that influence the educational process (such as habit formation) and design digital solutions for these key elements. By creating a ”perception-analysis-feedback” loop, we aim to achieve the digitization of early childhood education.
The purpose of this system is to provide more children with access to high-quality early childhood education opportunities and reduce the existing inequities. Simultaneously, we will leverage artificial intelligence and digital technologies to continuously improve the quality of education.
1.2 What We Will Do
In the following content, we will first analyze the application of AI in education, and further analyze the problems that exist in its practical application. Finally, we will propose our own solutions.
In addressing educational inequity, we focus on using internet technology and artificial intelligence to ensure that everyone has the same opportunity to receive high-quality education. In terms of digitizing the educational process, we will first define the key factors required for digital education and conduct research based on these factors. Throughout the process, multimodal human-computer interaction technologies, including affective computing, will be continuously applied to capture and influence the key elements of the digital education process.
We believe AI technology has the potential to help the world address these chronic challenges in education relating to equity, learning outcomes, and real world relevance.
2 Children’s Preschool Education and the Challenges It Faces
2.1 AI in Education
Baker et al. (2019) categorize AI in Education (AIED) applications into 3 categories: (i) learner-oriented AIED; (ii) instructor-oriented AIED; and (iii) institutional system-oriented AIED. However often applications that focus on learning cut cross several stakeholders especially in the current context and impact a common objective - learning. They are likely to reflect a collaboration among teachers, learners, parents, tutors and peers. We prefer the following modified categorization:
  • (i)
    learning-oriented, including learners, teachers, educational institutions, tutors, peers, parents and para educators;
  • (ii)
    institutional operations oriented financial aid, retention, learner experience;
  • (iii)
    policy oriented including governmental interventions, subsidies, competitive assessments like MCAS in the U.S.
Guan et al. (2020); Srinivasan (2022a) summarize the various definitions of AIED in prior research:
Authors Definition
Hwang (2003)
Intelligent tutoring system that helps to organize system knowledge and operational
information to enhance operator performance and automatically determining exercise
progression and remediation during a training session according to past student performance.
Johnson et al. (2009)
Artificially intelligent tutors that construct responses in real-time using its own
ability to understand the problem and assess student analyses.
Popenici and Kerr (2017)
Computing systems that are able to engage in human-like processes such as learning,
adapting, synthesizing, self-correction and use of data for complex processing tasks.
Chatterjee and Bhattacharjee (2020)
Computing systems capable of engaging in human-like processes such as adapting, learning,
synthesizing, correcting and using of various data required for processing complex tasks.
The first two definitions of AIED above are focused on intelligent tutoring and the last two are slightly broader but still focused on the learner.
Srinivasan (2022b) mentioned, based on an analysis of over 400 research articles published between 2000 and 2019 on the application of AI for teaching and learning, they find shifts in research emphasis between the first and second decade. During the period 2000–2009, the emphasis was on learner-oriented approaches trending away from instructor-oriented approaches. In the second decade 2010–2019, they found that there was a noticeable shift towards modeling learning outcomes/analytics and student profiling. Their findings reflect the growing interest in explicit measurement of learning outcomes, and predictive AI models including the use of deep statistical learning.
Based on a review of 220 research articles on personalized language learning, Chen et al. (2021) find that (1) multimodal videos promote effective personalized feedback; (2) personalized context-aware ubiquitous language learning enables active interaction with the real world by applying authentic and social knowledge to their surroundings; (3) mobile chatbots with Automated Speech Recognition (ASR) provide human-like interactive learning experiences to practice speaking and pronunciation; (4) collaborative game-based learning with customized gameplay path and interaction and communication features supports social interactivity and learning of various language skills; (5) AI promotes effective outcome prediction and instruction adaptation for individuals based on massive learner data; (6) Learning Analytics (LA) dashboards facilitate personalized recommendations through learning data visualization; and (7) data driven learning allows personalized and immediate feedback in real-time practice. Generally, these findings validate that AI technology can be beneficial in personalized language learning.
The above reviews provide valuable summaries of the progress AIED has made, especially in personalized and adaptive learning. Broadly the findings support the hypothesis that personalization makes a difference when learner characteristics and differences are carefully considered. The trend towards more explicit measurement of learning outcomes is healthy and much needed. In incorporating learning styles, extant research with the exception of Srinivasan and Murthy (2021) does not consider the beneficial impact of engaging multiple sensory channels. The extant research also does not explicitly propose a generalized framework for integrating findings from learning science. Granular personalization governed by learning styles including the influence of different sensory modes is integral to a more holistic AI-led learning framework.
2.2 Current Limitations
Analyzing the work of AI in education, we found that an effective AI learning system in early childhood education typically possesses the 7 characteristics mentioned by Chen et al. (2021). However, most existing work tends to focus on only one aspect of these characteristics, and there is a lack of a medium to integrate these 7 features into a cohesive educational system.
This medium must have characteristics that allow it to be generated or processed by each link and be relevant to the educational process environment. Most importantly, this medium must have a certain degree of manipulability, so that by manipulating it, the educational process can be influenced.
Through investigation and analysis, we found that research on the mechanisms of habit formation (LEWIS et al. (2021); Lewis et al. (2021)) may reveal an excellent medium suitable for AI in education, which is the factors influencing habit formation. These factors not only influence habit formation but are essentially key elements affecting learning behavior. Moreover, these factors can be quantified, observed, and altered, making them perfectly applicable to research on AI in preschool education.
I believe the key here is to understand how to extract the critical factors that might influence or drive educational behavior from the three interactive scenarios of ”child-teacher”, ”child-environment”, and ”teacher-environment”. By analyzing and optimizing these critical factors, we can enhance the success rate of educational behavior.
2.3 My Solution
The factors associated with preschool education can be easily extracted or applied in ”(1) multimodal videos”. AI in human identification based on these factors can be easily adopted for ”(2) personalized context-aware systems”. Large language models can be easily adopted in ”(3) mobile chatbots to extract important factors”. Multimodal interaction based on factor guidance or extraction can be applied to ”(4) collaborative game-based learning” and ”(5) AI for effective outcome prediction”. Graph knowledge-based algorithms can be easily applied to factor analysis in ”(6) Learning Analytics (LA)”. Factors can also be collected for ”(7) data-driven learning”.
Therefore, our solution will revolve around key factors, focusing on the extraction, identification, and feedback of critical factors affecting the educational process. Specifically, we will design the system to address the following three aspects.
2.3.1 How to Extract Factors Related to Study in Human-Computer Interaction?
The interaction between children, teacher, and their environment plays an important role in their education process. By changing the environment, children and teacher’s behaviors can also be altered. For example, when there are many unfamiliar adults present, children tend to behave more conservatively. Conversely, when the environment is filled with toys, children often feel more relaxed and exhibit more open behaviors. Additionally, certain internal states of children can also affect their behavior. For instance, a happy child is often more willing to engage in conversation.
Investigating how to identify key factors from the interaction between children, teacher, and their environment that may influence children’s behavior and ultimately their education process is a valuable research topic.
2.3.2 How to Identify Important factors?
In the past, researchers have been working to identify factors that are important for education (like habit formation), both environmental and individual. These factors include positive factors that can promote education, and negative factors that can inhibit education.
However, the important factors often vary among individuals, making it challenging to identify a single set of crucial elements for children’s preschool education. More importantly, data-driven deep learning algorithms need to handle large-scale data, and current simple models based on strategies such as random forests and linear regression often fail to meet the requirements. Therefore, algorithms capable of identifying important factors in large-scale data are crucial.
2.3.3 How to Use Important Factors to Guide Behavior?
After addressing the two issues mentioned above, we will encounter a new research topic: how to use the identified key factors to change children or teacher’s behavioral for higher educational success rate.
2.4 Summary
By addressing the three issues mentioned above, our system will form a closed loop.
Specifically, the system will interact with children and teachers to understand their behavioral habits and identify important factors that may improve education quality. Once these important factors are identified, the system will use them to further influence the behaviors of children and teachers, thereby contributing to the improvement of education quality, as illustrated in Figure 1.
Figure 1: The construction of an education system based on human-computer interaction that I envision. This system comprises three components. The first component is the use of artificial intelligence algorithms to extract education-related elements from interactions. The second component involves using artificial intelligence algorithms to identify factors that might have an impact. The final component employs reinforcement learning algorithms to regulate learning and educational behaviors, ultimately enhancing the quality of education.
3 Hardware and software system
Human-computer interaction is a highly effective method for extracting information and affective computing. In our interactive education system, the hardware system is primarily used for students or teachers to interact with the backend algorithms. The software system mainly processes these interaction data, while also providing feedback to the teaching process by adjusting the parameters of the hardware system.
3.1 Hardware Interaction System
To efficiently extract key factors in human-computer interaction, we first designed a hardware system for use in school or home settings, as shown in Figure 2.
Figure 2: The hardware framework of our multimodal interaction system (for demonstration purposes only, see Table LABEL:tab:hardware_list for the quantity and specifications of the equipment)
In our initial concept for the hardware system, it includes audio receivers, speakers, monitor cameras, computer, tablets, and laser projector, etc.
These devices can be used to capture interactions between children and teachers and perform related affective computing and educational behavior analysis. They can also directly interact with teachers or children, conduct data analysis, and provide status feedback.
In the initial phase of our research, we chose to use the free facilities provided by AIRS to set up the hardware system. Subsequent research will be based on this hardware. These free facilities can significantly reduce the cost of the hardware system and improve the efficiency of system setup.
3.2 Software Interaction System
Based on the aforementioned hardware system setup, we will also develop the corresponding software system.
In our software system, artificial intelligence algorithms play an important role. Our software system can provide rich multimodal interaction capabilities. With the support of AI algorithms, our software system can perform real-time analysis and collection of teaching data, as well as provide guidance and interventions in the educational process.
The overall software framework are illustrated in Figure 3. In our software system, in addition to developing some AI algorithms ourselves, we have also procured some mature technologies provided by third parties. For specific details, please refer to Table LABEL:tab:softwarecost.
Figure 3: Diagram of the proposed multimodal human-computer interaction software system framework
This software system models the digital education process around multimodal input through methods such as affective computing, computer vision, machine learning, and data analysis.
Additionally, we use large model technology to provide realistic dialogue interactions, enhancing the teaching process’s fun by offering children anthropomorphic interaction methods, thereby improving the quality of interaction and data.
This system is designed around the extraction, analysis, and feedback of factors during the interaction process, forming a complete digital education system. In the following content, we will focus on two factor extraction strategies under different interaction methods, including affective computing and behavior and emotion analysis based on LLM (Large Language Models).
4 Extracting Factors Related to Education in Human-Computer Interaction (HCI)
Before analyzing the educational process, we need to identify the key factors that influence it, including children’s emotions, knowledge retention, teachers’ emotions, teaching methods, and environmental temperature, among others.
To achieve this goal, we first need to design a user-friendly multimodal interaction system. The design principles of these systems should ensure that they are appealing to children, teachers, and other participants. van der Meer et al. (2012) and McLay et al. (2017) pointed that children with autism spectrum disorders (ASD) show a preference for one of the three communication modes: manual signing, picture exchange, and speech-generating devices. In addition, Anderson et al. (1994) pointed that successful communication involves active involvement of both participants, asking and answering questions, volunteering information, and responding sensitively to contributions from their partners.
Based on these research findings, we discovered that an interactive system appealing to children must first exhibit behavior patterns similar to those of humans. This includes voice conversations, asking questions, answering questions, providing additional information, and responding to children’s emotional changes.
Based on our designed multimodal interaction system of hardware and software, this section introduces several possible methods for interactive factor extraction.
4.1 Extraction of Key Factors Based on Affective Computing
Emotion is generally considered an important factor influencing education, and this factor can be conveniently implemented using facial recognition.
The research content of affective computing primarily involves the following five aspects:
  • Fundamental Theory
  • Emotion Signal Collection
  • Multimodal Integration
  • Emotion Generation and Expression
  • Emotion Analysis
In fundamental theoretical research, emotion models are divided into two types: discrete emotion models and dimensional emotion models. In our research plan, we use the dimensional emotion model for emotion analysis because the data in this modality is easier to quantify and can be more readily processed by AI algorithms.
In the current field of affective computing, data from sensors such as cameras and microphones are primarily used. Additionally, physiological data, which straightforwardly and objectively reflect an individual’s state, have garnered increasing attention from researchers in recent years. Currently, commonly used physiological data include electroencephalogram (EEG), electrodermal activity (EDA), respiration, skin temperature, electrocardiogram (ECG), electromyogram (EMG), blood volume pulse (BVP), and electrooculogram (EOG). However, due to the impact of cost and the manner of wearing the sensors, the amount of such data is relatively limited in current research.
4.1.1 Affective Computing based on Speech Data
Figure 4: Emotional computing based on speech data (Hu et al. (2019))
Based on voice emotion analysis, current research is very well-developed. These methods leverage relevant techniques in linguistics and acoustics, not only analyzing grammar and semantics but also identifying acoustic features related to emotional states, such as speech rate and intonation. Currently, the most widely used tools for extracting emotional speech features are openSMILE and Librosa. In our system, we will directly use Librosa’s technology to extract emotional features from speech.
A typical example is shown in Figure 4.
4.1.2 Affective Computing based on Videos and Images
In emotion recognition from videos and images involving facial expressions, body movements, and environmental contexts, the recognition of facial expressions constitutes the main focus of research.
The relevant technological developments are already very mature, as can be seen in the work of Li et al. (2020). Therefore, we will directly use these mature technologies for the analysis and collection of emotional factors. An example of emotion analysis in a typical educational scenario is shown in Figure 5.
Figure 5: Typical applications based on facial emotion analysis
4.1.3 Affective Computing based on Video and Dialogue Multimodal Fusion
The multimodal fusion of text and video is very important when extracting factors, as it can provide multi-scale perspectives to more accurately extract factors from the teaching process.
To better extract factors, we use Vision LLM to converse with children, or use conversations between teachers and children, along with video information during the process to extract key factors. The structure is shown in Figure 6.
Figure 6: Important factors extraction based on video and conversation
In this process, we can see that the Vision LLM-based multimodal algorithm can capture the corresponding states of children in various daily life scenarios and record these states in the form of dialogues. Ultimately, it obtains the corresponding factors through structured output. Specifically, in our software system, we use cameras to capture children’s facial expressions and engage in voice conversations with them, asking their opinions about the current task. Through multiple rounds of dialogue, we gather the essential features of children’s perceptions of specific subjects.
Vision Large Language Model (V-LLM) is fully capable of performing this task. Because the V-LLM can communicate with children like a human and always remains patient. Additionally, the V-LLM provides emotional value feedback to children through storytelling and other methods, thereby attracting children to engage more willingly in communication.
It is worth noting that the children’s voices will be converted into text and input into Vision LLM, while the responses generated by the Vision LLM will be converted into audio and played for the children to hear.
Compared to simply recognizing facial expressions, this approach can provide individuals with a richer set of optional factors, thereby ensuring high-quality importance results.
4.1.4 Physiological Data Analysis
Electroencephalogram (EEG) data has been widely studied in the field of physiological data analysis. Since 2018, there has been a significant increase in the number of papers utilizing deep learning methods for emotion computation using EEG data. These methods include Convolutional Neural Networks (CNN), Deep Belief Networks (DBN), Recurrent Neural Networks (RNN), and Stacked Auto Encoders (SAE), among others.
We will primarily handle data in modalities such as images and speech initially, and this type of data may be further studied in the future.
4.1.5 A Multimodal Coding System for Student Behavior
Multimodal sentiment analysis can lead to more accurate results. Common multimodal fusion techniques include feature-based fusion, model-based fusion, and decision-based late fusion. In future research, we will explore these three aspects according to the requirements of specific tasks.
The current form of the multimodal encoding system for smart education is shown in Table LABEL:tab:coding.
Table 1: Multimodal data coding system.
Object Enviroment Multimodal data Learning Behavior Coding
Qualitative
characterization
Student Intelligent recording system Sound data Audio Data Valid Active response S1 Proactive
Passive response S2 Participation
Ask questions
proactively
S3 Proactive
Discussion S4 Proactive
Sharing of results S5 Proactive
Evaluation feedback S6 Proactive
Invalid
Ineffective confusion
(topic deviation,
intentional confusion)
S7 Participation
Silent Data Valid Think S8 Proactive
Listening S9 Proactive
Invalid Ineffective Silence S10 Participation
Intelligent recording and broadcasting system image data Location Move Valid Seat M1 Participation
Podium M2 Participation
Between groups M3 Participation
In group M4 Participation
Invalid Corridor M5 Participation
No location
movement
No location
movement
M6 Participation
Body Movements Valid
Look at the
interactive whiteboard
A1 Focus
Watch the teacher A2 Focus
Watch classmates A3 Focus
Watch the
learning tools
A4 Focus
Looking at
the book
A5 Focus
For interactive
whiteboard
A6 Participation
For teachers A7 Participation
For companions A8 Participation
For learning tools A9 Participation
Book-oriented A10 Participation
Invalid Not related to learning A11 Participation
No body movements Sleep A12 Participation
Distracted A13 Participation
Learning platform (Wechat,MOOC etc.) text data Text data used by technology Valid technical support Sign in T1 Participation
Answer T2 Participation
Homework Exercises T3 Participation
Self-learning T4 Participation
Discussion T5 Participation
Sharing of results T6 Participation
Evaluation feedback T7 Participation
Data collection T8 Participation
Invalid technical support
Unrelated to
learning technology use
T9 Participation
No technology usage T10 Participation
This will serve as a reference in the early stages of project development.
4.2 Storing Various Factors for Different Children in a Graph Database
To effectively organize the data, we extract key factors for each child individually and store these key factors in a graph database, as shown in Figure 7.
Figure 7: Diagram of Storing Data Using a Graph Database
In this data organization structure, we decouple time, attributes, and different children, making it more suitable for large-scale computation and processing. Additionally, this approach provides more opportunities to explore how to use knowledge graph-related technologies to investigate the deep relationships between habit formation in children and various factors.
5 Identifying Important Factors in Education Process
The habit formation is somewhat related to the educational process. For example, both have a goal, and both are influenced by certain key factors. Initially inspired, we found that we can use algorithms that identify important factors relevant to the field of habit formation to pinpoint factors that significantly impact education.
We first introduce a classic method for calculating key factors based on the Growth Curve Model (Lewis et al. (2021)) and Shapley value (LEWIS et al. (2021)). Then, we propose a new algorithm based on Knowledge Graph methodologies to address the issues faced by these two methods.
5.1 Growth Curve Model in Important Factors Identification
Assuming we have 5 participants who record their habit automaticity scores (SRHI Automaticity) every day for 21 days. Additionally, we collect each participant’s initial commitment, previous mindfulness experience, and personality traits, including agreeableness and emotional stability.
The collected sample format is illustrated in Table 2.
Table 2: Table caption. For brevity, only a subset of the data is shown here
Participant Day SRHI Automaticity Commitment Experience Agreeableness Emotional Stability
1 1 2.0 High Low High Low
1 2 2.2 High Low High Low
1 21 3.5 High Low High Low
2 1 1.5 Medium Medium Low High
2 2 1.7 Medium Medium Low High
5 21 4.2 Low High Medium Medium
First, we aggregate the daily data of all participants into a long-format data frame, where each row represents a participant’s record on a specific day.
Then we use a Growth Curve Model to analyze the data. The hierarchical structure of the model is as follows:
  • Level-1 (Time Variation): Describes the change in individual habit automaticity over time.
  • Level-2 (Individual Differences): Describes the differences between individuals, using time-invariant covariates such as initial commitment, experience, agreeableness, and emotional stability.
The model equations can be represented as follows:
$y_{ti}^{SRHI}A=b_{1i}+b_{2i}\cdot t+u_{ti}$ (1)
where $b_{1i}=\beta_{01}+\beta_{11}\cdot Commitment_{i}+\beta_{21}\cdot Experience_{i}+\beta_{31}\cdot Agreeableness_{i}+\beta_{41}\cdot EmotionalStability_{i}+d_{1i}$, $b_{2i}=\beta_{02}+\beta_{12}\cdot Commitment_{i}+\beta_{22}\cdot Experience_{i}+\beta_{32}\cdot Agreeableness_{i}+\beta_{42}\cdot EmotionalStability_{i}+d_{2i}$.
After obtaining the model equations, we can use Python to fit the above model and compute the corresponding parameters. The code is as follows:
1import pandas as pd
2import numpy as np
3import statsmodels.api as sm
4from statsmodels.formula.api import mixedlm
5
6np.random.seed(42)
7days = np.arange(1, 22)
8data = []
9
10for participant in range(1, 6):
11 commitment = np.random.choice([’High’, ’Medium’, ’Low’])
12 experience = np.random.choice([’High’, ’Medium’, ’Low’])
13 agreeableness = np.random.choice([’High’, ’Medium’, ’Low’])
14 emotional_stability = np.random.choice([’High’, ’Medium’, ’Low’])
15
16 for day in days:
17 # Generate random SRHI Automaticity data (sample data)
18 srhi_automaticity = np.random.normal(loc=3 + 0.1 * day, scale=0.5)
19 data.append([participant, day, srhi_automaticity, commitment, experience, agreeableness, emotional_stability])
20
21df = pd.DataFrame(data, columns=[’participant’, ’day’, ’srhi_automaticity’, ’commitment’, ’experience’, ’agreeableness’, ’emotional_stability’])
22
23# Convert categorical variables to dummy variables
24df = pd.get_dummies(df, columns=[’commitment’, ’experience’, ’agreeableness’, ’emotional_stability’], drop_first=True)
25
26# Define formula
27formula = ’srhi_automaticity ~ day + commitment_Medium + commitment_Low + experience_Medium + experience_Low + agreeableness_Medium + agreeableness_Low + emotional_stability_Medium + emotional_stability_Low’
28
29# Fit a mixed-effects linear model
30model = mixedlm(formula, df, groups=df[’participant’], re_formula=’~day’)
31result = model.fit()
32
33# Output the model results summary
34print(result.summary())
After fitting the model, you can review the model parameters and significance levels. The fitted results are typically in the following form:
1Mixed Linear Model Regression Results
2=============================================================
3Model: MixedLM Dependent Variable: srhi_automaticity
4No. Observations: 105 Method: REML
5No. Groups: 5 Scale: 0.2765
6Min. group size: 21 Log-Likelihood: -95.2805
7Max. group size: 21 Converged: Yes
8Mean group size: 21.0
9-------------------------------------------------------------
10 Coef. Std.Err. z P>|z| [0.025 0.975]
11-------------------------------------------------------------
12Intercept 2.710 0.130 20.846 0.000 2.455 2.965
13day 0.010 0.005 2.102 0.036 0.001 0.020
14commitment_Medium -0.030 0.100 -0.300 0.764 -0.226 0.166
15commitment_Low 0.100 0.120 0.833 0.405 -0.135 0.335
16experience_Medium 0.030 0.010 3.000 0.003 0.010 0.050
17experience_Low -0.030 0.020 -1.500 0.134 -0.070 0.010
18agreeableness_Medium -0.020 0.010 -2.000 0.045 -0.040 0.000
19agreeableness_Low 0.020 0.010 1.800 0.072 -0.002 0.042
The new question is, how can we determine which factors have a significant impact on habit formation from this result?
The answer lies in observing the regression coefficients and significance levels. Specifically, in the model results summary, we focus on each covariate’s regression coefficient (Coef.), standard error (Std. Err.), z-value (z), and p-value (P $>|z|$). The p-value is used to determine the significance of each covariate’s impact on the change in habit automaticity. Typically, we use the following criteria to judge significance:
  • p $<$ 0.05: The covariate has a significant effect on the outcome variable.
  • p $<$ 0.01: The covariate has a highly significant effect on the outcome variable.
  • p $<$ 0.1: The covariate has a marginally significant effect on the outcome variable.
For significant covariates, the regression coefficient indicates the direction and magnitude of that variable’s effect on habit automaticity scores. For example:
  • A positive regression coefficient indicates that an increase in the covariate leads to an increase in habit automaticity scores.
  • A negative regression coefficient indicates that an increase in the covariate leads to a decrease in habit automaticity scores.
Therefore, we can analyze the model output as follows:
  • Time Effect (day):
    Regression coefficient is 0.01, p-value is 0.036 (p ¡ 0.05), indicating that habit automaticity significantly increases over time.
  • Commitment:
    commitment Medium: Regression coefficient is -0.03, p-value is 0.764 (p ¿ 0.05), indicating that medium commitment has no significant effect on the change in habit automaticity.
    commitment Low: Regression coefficient is 0.10, p-value is 0.405 (p ¿ 0.05), indicating that low commitment has no significant effect on the change in habit automaticity.
  • Mindfulness Experience (Experience):
    experience Medium : Regression coefficient is 0.03, p-value is 0.003 (p ¡ 0.01), indicating that medium mindfulness experience has a highly significant positive effect on the increase in habit automaticity.
    experience Low: Regression coefficient is -0.03, p-value is 0.134 (p ¿ 0.05), indicating that low mindfulness experience has no significant effect on habit automaticity.
  • Agreeableness:
    ‘agreeableness Medium: Regression coefficient is -0.02, p-value is 0.045 (p ¡ 0.05), indicating that medium agreeableness has a significant negative effect on habit automaticity.
    agreeableness Low: Regression coefficient is 0.02, p-value is 0.072 (p ¡ 0.1), indicating that low agreeableness has a marginally significant positive effect on habit automaticity.
  • Emotional Stability:
    emotional stability Medium‘: Regression coefficient is 0.02, p-value is 0.072 (p ¡ 0.1), indicating that medium emotional stability has a marginally significant positive effect on habit automaticity.
    emotional stability Low: Regression coefficient is 0.02, p-value is 0.072 (p ¡ 0.1), indicating that low emotional stability has a marginally significant positive effect on habit automaticity.
Thus, we can conclude that the factors significantly affecting habit formation are as follows:
  • Time Effect (day): Habit automaticity significantly increases over time.
  • Mindfulness Experience (Experience): Medium mindfulness experience significantly promotes an increase in habit automaticity.
  • Agreeableness: Medium agreeableness significantly inhibits the increase in habit automaticity.
  • Emotional Stability: Both medium and low emotional stability have a marginally significant positive effect on habit automaticity.
Through this analytical approach, we can identify which individual characteristics significantly influence the speed of habit formation, thereby providing a basis for personalized interventions.
5.2 Shapely Algorithm in Important Factors Identification
LEWIS et al. (2021) using the Shapley algorithm (Winter (2002)), important factors influencing habit formation are extracted from survey data. This process mainly includes the following steps:
  • Data Collection: Collect data from participants’ daily surveys, including whether they practiced breathing exercises, the level of automation, sense of reward, sense of challenge, confidence, motivation, mood, and daily activities.
  • Model Training: Use the collected data to train a binary classification model (e.g., Random Forest) to predict whether participants will practice breathing exercises the next day.
  • Shapley Value Calculation: For each model prediction, calculate the Shapley value for each feature, measuring its contribution to the prediction result. Shapley values represent the marginal contribution of features to the prediction result.
  • Feature Importance Analysis: By aggregating Shapley values, determine which features are most important for model predictions.
In summary, the Shapley method is more direct compared to the Growth Curve Model. The Growth Curve Model requires calculating significant statistical information from long-term surveys to determine the important factors influencing habit formation. In contrast, Shapley can calculate this importance by processing a single day’s data, and further improves prediction accuracy by accumulating data over multiple days.
Below is a simplified Python implementation process, demonstrating how to use the Shapley value algorithm to extract important features from constructed data.
First, we need to construct a simulated dataset. Let’s assume we have the following features:
  • did breathing: Whether the breathing exercise was performed (0 or 1)
  • automaticity: Level of automation (rating from 1 to 7)
  • rewarding: Sense of reward (rating from 1 to 7)
  • challenging: Sense of challenge (rating from 1 to 7)
  • confidence: Confidence (rating from 1 to 7)
  • motivation: Motivation (rating from 1 to 7)
  • mood: Mood (rating from 1 to 7)
  • busy: Level of busyness (rating from 1 to 7)
These data can be implemented and processed using the following Python code:
1import numpy as np
2import pandas as pd
3from sklearn.ensemble import RandomForestClassifier
4from sklearn.model_selection import train_test_split
5import shap
6
7np.random.seed(42)
8data_size = 1000
9data = pd.DataFrame({
10 ’did_breathing’: np.random.randint(0, 2, data_size),
11 ’automaticity’: np.random.randint(1, 8, data_size),
12 ’rewarding’: np.random.randint(1, 8, data_size),
13 ’challenging’: np.random.randint(1, 8, data_size),
14 ’confidence’: np.random.randint(1, 8, data_size),
15 ’motivation’: np.random.randint(1, 8, data_size),
16 ’mood’: np.random.randint(1, 8, data_size),
17 ’busy’: np.random.randint(1, 8, data_size)
18})
19
20# Features and Labels
21X = data.drop(’did_breathing’, axis=1)
22y = data[’did_breathing’]
23
24# Traning and Validation datasets
25X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2, random_state=42)
26
27# Train Random Forest Classifier
28model = RandomForestClassifier(n_estimators=100, random_state=42)
29model.fit(X_train, y_train)
30
31# Using SHAP to Explain the Model
32explainer = shap.TreeExplainer(model)
33shap_values = explainer.shap_values(X_test)
34
35# Features Importance Visualization
36shap.summary_plot(shap_values[1], X_test)
37
38# Print Absolute Averaged Shapley Value
39shap_importance = np.abs(shap_values[1]).mean(axis=0)
40feature_importance = pd.DataFrame(
41 list(zip(X_train.columns, shap_importance)),
42 columns=[’Feature’, ’Importance’]
43).sort_values(by=’Importance’, ascending=False)
44
45print(feature_importance)
Through the above process, we can determine which features are most important to the model’s predictions, thereby understanding which factors significantly influence habit formation.
5.3 Knowledge Graph Algorithms in Important Factors Identification
The Growth Curve Model and Shapley value method both employ very simple linear fitting or simple classification models, which are very effective when dealing with small amounts of data (e.g., 62 participants). However, such algorithms cannot handle large-scale data. Additionally, these algorithms do not consider the differences between individuals; important factors are directly derived from the averages of statistical data, assuming these factors are important for everyone. This is intuitively unreasonable.
In contrast, Graph Neural Networks (GNNs) can be used to handle complex relationships and structured data, enabling them to capture relationships within the data better than traditional machine learning models in some cases.
To better distinguish individual differences and conduct habit formation research on large-scale data, we first store all children and related factors in a graph database (we chose Neo4J, which can be easily interacted with using Python). Specific details can be found in Chapter 2.3.
Inspired by the work of LEWIS et al. (2021), we use Graph Neural Networks (such as Graph Convolutional Networks, GCN (Zhang et al. (2019)), or Graph Attention Networks, GAT (Veličković et al. (2017))) to train the model. The model can propagate information through the graph via a message-passing mechanism, capturing more complex relationships. The goal of the GNN model is to predict whether each participant will perform a breathing exercise next time.
Although GNNs are inherently black-box models, techniques similar to Shapley values (such as GNNExplainer Ying et al. (2019)) can be used to explain the contribution of each node and edge to the final prediction. GNNExplainer optimizes a subgraph to maximize the explanation of the model’s prediction, thereby identifying which nodes and edges are most important for the prediction.
6 Regulating Children’s Behavior Using Important Factors
To regulate children’s behavior using important factors, we first need to understand the basic principles of behavior control (Frings et al. (2020)). Traditional theories suggest that behavior is controlled by two distinct systems: the goal-directed system and the habit system.
The goal-directed system refers to the system that makes decisions and takes actions based on the consequences of behavior. This system requires individuals to have a clear understanding of the relationship between behavior and its outcomes to make appropriate choices. Goal-directed behavior primarily involves the brain’s prefrontal cortex and dorsomedial striatum (DMS). These areas are responsible for integrating information, planning, and decision-making. Therefore, the goal-directed system is more suited to using reinforcement learning. One classic experiment is the reward devaluation experiment, where animals reduce behaviors related to rewards that are no longer attractive (for example, when food is poisoned).
The habit system refers to the behavior system triggered automatically by specific situations or stimuli, rather than controlled by the consequences of behavior. This system is formed through repetition and reinforcement, making behavior automatic and not requiring consideration of consequences. Habit behavior primarily involves the brain’s dorsolateral striatum (DLS). These areas are related to motor control and automated behaviors. The habit system is more like a stress response, such as stopping at a red light every time you see it.
The goal-directed system and the habit system represent two extreme forms of behavior control. The goal-directed system allows individuals to flexibly adjust their behavior to achieve optimal outcomes, while the habit system automates behavior, thereby conserving cognitive resources. In real life, most behaviors result from the combined action of these two systems, dynamically adjusting between them based on context and experience.
In recent years, reinforcement learning has gradually gained attention and has been introduced into the study of behavior control (Fan et al. (2023); Wu and Wang (2023)). In the following content, I will discuss in detail the current theoretical models of behavior control and, based on these models, propose my own framework for regulating children’s behavior by incorporating important factors.
6.1 Binding and Retrieval in Action Control (BRAC)
6.1.1 Basic Concepts
Frings et al. (2020) believe that human behavior control relies on representations that integrate perception and action. By associating specific behaviors with their outcomes, humans learn to act purposefully. Based on this cognitive hypothesis, Frings et al. (2020) proposed the BRAC framework, aimed at integrating and explaining a wide array of research findings on human behavior control.
Specifically, the BRAC framework proposes two core processes: feature binding and retrieval. Binding refers to integrating the features of stimuli, responses, and effects into an event file; retrieval refers to automatically retrieving previously bound event files when a particular feature is repeated. The framework emphasizes that binding and retrieval are functionally distinct and separable processes, and these two processes are independently regulated by top-down and bottom-up processes.
The BRAC framework can integrate and explain phenomena of behavior control across various experimental paradigms, such as task switching, negative priming, and repetition priming. Through this framework, different behavior control tasks can be reinterpreted as outcomes driven by binding and retrieval processes.
In addition, BRAC can also be extended to other related fields such as attention, memory and learning, and motivation. For example, in attention research, feature binding and retrieval can explain cross-trial effects in spatial orientation and visual search tasks; in memory and learning research, the concept of event files is closely related to episodic memory; in motivation research, BRAC can explain how approach and avoidance behaviors influence the evaluation of neutral stimuli.
The BRAC framework helps to unify and simplify theoretical explanations of behavior control phenomena, facilitating the integration of results across different experimental paradigms. The framework also raises new hypotheses and research questions, such as how to distinguish the specific contributions of binding and retrieval processes in behavior control, and what the neural bases of these processes are.
6.1.2 Children’s Behavior Regularization based on Important Factors and BRAC
The BRAC theory emphasizes two core processes in behavior control: feature binding and retrieval. Inspired by this theory, we can utilize it to influence behavior by changing the processes of feature binding and retrieval of important factors.
Specifically, we can design behavior interventions in three aspects: enhancing positive feature binding, reducing negative feature binding, and increasing the effectiveness of feature retrieval.
In enhancing positive feature binding, we can strengthen the positive feature binding associated with target behavior through positive experiences and feedback. For example, we can design reward mechanisms that offer incentives for completing habit practices, such as points, badges, or small gifts, to enhance the sense of reward from practice. Additionally, we can use LLMs to provide instant positive feedback, such as encouraging words or progress charts, to increase the enjoyment and sense of accomplishment from practice.
In reducing negative feature binding, we can minimize negative experiences and emotional bindings associated with target habits. For instance, we can lower the difficulty of habit practice by breaking tasks down or providing guidance, making them easier to complete. We can also introduce relaxation techniques or time management strategies to help individuals better cope with stress and busy schedules.
In increasing the effectiveness of feature retrieval, we can enhance the retrieval process of relevant features through environmental cues and behavior prompts, promoting the automation of habits. For example, we can set reminders at key times and locations, such as phone notifications or calendar alerts, to remind individuals to practice their habits. We can also bind target habits to existing daily behaviors, such as meditating after brushing teeth, leveraging the automation of existing habits to trigger new ones.
In addition to these interventions, we can design personalized interventions based on the characteristics and needs of different individuals. For example, adjusting the intensity and content of interventions based on different emotional states and levels of busyness. Specifically, during low emotional states, we can try increasing positive feedback and rewards, providing more emotional support. During busy states, we can simplify the steps of habit practice and offer time management suggestions.
By continuously collecting individual feedback and behavior data, we can dynamically adjust interventions to ensure their continued effectiveness. For instance, using smartphone applications to track individuals’ habit practice status and adjusting reminders and feedback strategies based on the data.
6.2 Optimal Behavior Estimation by Reforcement Learning
BRAC has conducted systematic research on the role of stress responses in individual behavior, but it has yet to address an important issue: what specific actions should an individual take when they are in a certain state? For instance, when a person is sad, they might need to talk to someone or silently endure. This decision depends on the individual’s factors, such as personality.
However, how to use a behavior intervention system to help individuals overcome negative emotions and make optimal behavioral decisions is still a question worth exploring. Furthermore, how to utilize reinforcement learning to determine the optimal behavior, thereby achieving positive goal (such as efficient habit formation) is an area worthy of research.
To this end, a feasible solution is to first use reinforcement learning to build a behavior model. The output of this model would be the action the user should take in their current state, and this action would always tend to obtain more rewards (e.g., it would more easily help form good habits). The overall framework of this system is shown in Figure 8.
Figure 8: Motivation for Optimal Behavior Estimation Based on RL
In this system, we will use the techniques from Chapters 2 and 3 to extract important factors as state inputs for the reinforcement learning model. For a preference-trained reinforcement learning model, the output will be more inclined to produce actions that align with the user’s preferences (favoring the pursuit of a better emotional state).
6.2.1 Behavioral Network based on Reforcement Learning
Yamada and Toda (2023) proposed a Q-learning method that can effectively help us understand how to use reinforcement learning to model the learning of behaviors and the corresponding decision-making processes. Reinforcement learning is used to simulate the habit formation process within a behavior network.
This study assumes that behavior is a network composed of multiple responses, with reinforcement learning used to represent the agent’s history of rewards obtained during transitions between different responses. The agent uses the Q-learning algorithm to learn the Q-values (i.e., action values) between various responses, thereby guiding behavior selection and path searching.
The agent updates the Q-values using the Q-learning algorithm, where $Q(a_{t-1},a_{t})$ represents the action value of transitioning from response $a_{t-1}$ to response $a_{t}$. The Q-value update formula is:
$Q(a_{t-1},a_{t})\leftarrow Q(a_{t-1},a_{t})+\alpha\delta$ (2)
where $\alpha$ is the learning rate and $\delta$ is the reward prediction error. The formula for calculating the reward prediction error $\delta$ is:
$\delta=R(a_{t})+\gamma\max_{a_{t+1}}Q(a_{t},a_{t+1})-Q(a_{t-1},a_{t})$ (3)
where $R(a_{t})$ is the reward obtained when transitioning to response $a_{t}$, and $\gamma$ is the discount factor for future rewards.
The agent selects responses based on Q-values through a softmax function, which increases the probability of selecting responses with higher Q-values. The agent also uses shortest path search algorithms (such as Dijkstra’s algorithm) to find the shortest path from the current response to the target response in the behavior network, thereby optimizing the behavior sequence.
By simulating different training amounts, reward schedules, and selection environments, the study demonstrated how the reinforcement learning model can reproduce the habit formation phenomena observed in experiments. The simulation results indicate that habit formation can be viewed as a change in the structure of the behavior network rather than behavior controlled by two separate systems.
This work helps researchers establish a unified theoretical framework to explain the transition process between goal-directed behavior and habitual behavior.
6.2.2 Preference-Aligned Reinforcement Learning Algorithm
Inspired by the work of Yamada and Toda (2023) and others, we propose a preference-aligned reinforcement learning algorithm. As illustrated in Figure 9.
Figure 9: Preference-Aligned Reinforcement Learning Algorithm
First, a standard supervised learning method is used to train an initial model. Typically, this initial model is trained on a large-scale dataset to ensure it has basic task capabilities. Then, in our designed multimodal human-computer interaction system, the model is further refined using human feedback to make its outputs more inclined to produce positive results. Specifically, the model generates multiple candidate outputs for a given input. In our interactive system, human evaluators rank or score these candidate outputs, and the optimal output is selected based on preferences that lead to positive outcomes.
A reward model is trained using the human feedback data. This reward model learns how to evaluate the quality of the model outputs based on human preferences.
Furthermore, the trained reward model is used as a reward signal to further optimize the initial model through reinforcement learning algorithms (e.g., policy gradient methods). Common steps include using feedback from the reward model to update the policy, making the model more inclined to generate high-scoring outputs, and continuously evaluating the policy’s performance and optimizing it based on the feedback.
This process can be iterated, collecting feedback, training the reward model, and updating the policy in multiple rounds to continuously enhance the model’s performance and alignment.
7 Conclusion
In our implementation plan, the multimodal interaction-based system is divided into the following three parts:
  • Dynamic factor extraction based on artificial intelligence algorithms
  • Identification of key factors
  • Child behavior regulation system
We intend to build a comprehensive and closed-loop preschool education interaction system from these three aspects. The required technologies and their respective completion statuses are shown in the table below.
Table 3: The technical points required to implement the current system and their progress
Required Item Application
Current
Status
Factor Discovery System
Smart
Home
Hardware
Collect multimodal interaction information
of children in a fixed room
0%
Software Interaction Child identification
To identify the identity of the child currently
interacting
80%
Emotion Analysis Extracting emotional factors 80%
Digital Human
Communicate with the child, provide emotional
value, and extract structured factor information
90%
Entertainment System Improve user stickiness 0%
HCI Plugins
Extracting children’s behavioral factors in
third-party software
0%
Graph Storage
Using graph databases to store
structured information
100%
Important Factor Identification Growth Curve Data Collection Enabled by Factor Discovery System 50%
Analysis Method
Methods for analyzing data and
identifying important factors
100%
Shapely Data Collection Enabled by Factor Discovery System 50%
Analysis Method
Methods for analyzing data and
identifying important factors
100%
Knowledge Graph Data Collection Enabled by Factor Discovery System 50%
Analysis Method
Methods for analyzing data and
identifying important factors
60%
Child Behavior Regulation System
Behavior Regularization
based on
Important Factors and BRAC
Regulating children’s behavior 0%
Optimal Behavior Estimation by RL
RL in
behavior modeling
Behavioral Model 80%
PARL
Preference-aligned reinforcement
learning algorithm
10%
Based on the current progress, we predict that such a system can be implemented within two years, and the required personnel will not exceed two people (including myself).
The contributions of the proposed solution are as follows.
  • Proposed a complete multimodal interactive education system.
  • Divided the application of AI in the educational process into three parts and accordingly proposed innovative algorithms and solutions.

### v2

@@
 By analyzing and optimizing these critical factors, we can enhance the success rate of educational behavior.
-2.3 My Solution
-The factors associated with preschool education can be easily extracted or applied in ”(1) multimodal videos”.
-AI in human identification based on these factors can be easily adopted for ”(2) personalized context-aware systems”.
-Large language models can be easily adopted in ”(3) mobile chatbots to extract important factors”.
-Multimodal interaction based on factor guidance or extraction can be applied to ”(4) collaborative game-based learning” and ”(5) AI for effective outcome prediction”.
-Graph knowledge-based algorithms can be easily applied to factor analysis in ”(6) Learning Analytics (LA)”.
-Factors can also be collected for ”(7) data-driven learning”.
-Therefore, our solution will revolve around key factors, focusing on the extraction, identification, and feedback of critical factors affecting the educational process.
-Specifically, we will design the system to address the following three aspects.
-2.3.1 How to Extract Factors Related to Study in Human-Computer Interaction?
-The interaction between children, teacher, and their environment plays an important role in their education process.
-By changing the environment, children and teacher’s behaviors can also be altered.
-For example, when there are many unfamiliar adults present, children tend to behave more conservatively.
-Conversely, when the environment is filled with toys, children often feel more relaxed and exhibit more open behaviors.
-Additionally, certain internal states of children can also affect their behavior.
-For instance, a happy child is often more willing to engage in conversation.
-Investigating how to identify key factors from the interaction between children, teacher, and their environment that may influence children’s behavior and ultimately their education process is a valuable research topic.
-2.3.2 How to Identify Important factors?
-In the past, researchers have been working to identify factors that are important for education (like habit formation), both environmental and individual.
-These factors include positive factors that can promote education, and negative factors that can inhibit education.
-However, the important factors often vary among individuals, making it challenging to identify a single set of crucial elements for children’s preschool education.
-More importantly, data-driven deep learning algorithms need to handle large-scale data, and current simple models based on strategies such as random forests and linear regression often fail to meet the requirements.
-Therefore, algorithms capable of identifying important factors in large-scale data are crucial.
-2.3.3 How to Use Important Factors to Guide Behavior?
-After addressing the two issues mentioned above, we will encounter a new research topic: how to use the identified key factors to change children or teacher’s behavioral for higher educational success rate.
-2.4 Summary
-By addressing the three issues mentioned above, our system will form a closed loop.
-Specifically, the system will interact with children and teachers to understand their behavioral habits and identify important factors that may improve education quality.
-Once these important factors are identified, the system will use them to further influence the behaviors of children and teachers, thereby contributing to the improvement of education quality, as illustrated in Figure 1.
-Figure 1: The construction of an education system based on human-computer interaction that I envision.
-This system comprises three components.
-The first component is the use of artificial intelligence algorithms to extract education-related elements from interactions.
-The second component involves using artificial intelligence algorithms to identify factors that might have an impact.
-The final component employs reinforcement learning algorithms to regulate learning and educational behaviors, ultimately enhancing the quality of education.
 3 Hardware and software system
@@
 This system is designed around the extraction, analysis, and feedback of factors during the interaction process, forming a complete digital education system.
-In the following content, we will focus on two factor extraction strategies under different interaction methods, including affective computing and behavior and emotion analysis based on LLM (Large Language Models).
 4 Extracting Factors Related to Education in Human-Computer Interaction (HCI)
@@
 However, due to the impact of cost and the manner of wearing the sensors, the amount of such data is relatively limited in current research.
-4.1.1 Affective Computing based on Speech Data
-Figure 4: Emotional computing based on speech data (Hu et al. (2019))
-Based on voice emotion analysis, current research is very well-developed.
-These methods leverage relevant techniques in linguistics and acoustics, not only analyzing grammar and semantics but also identifying acoustic features related to emotional states, such as speech rate and intonation.
-Currently, the most widely used tools for extracting emotional speech features are openSMILE and Librosa.
-In our system, we will directly use Librosa’s technology to extract emotional features from speech.
-A typical example is shown in Figure 4.
-4.1.2 Affective Computing based on Videos and Images
-In emotion recognition from videos and images involving facial expressions, body movements, and environmental contexts, the recognition of facial expressions constitutes the main focus of research.
-The relevant technological developments are already very mature, as can be seen in the work of Li et al.
-(2020).
-Therefore, we will directly use these mature technologies for the analysis and collection of emotional factors.
-An example of emotion analysis in a typical educational scenario is shown in Figure 5.
-Figure 5: Typical applications based on facial emotion analysis
-4.1.3 Affective Computing based on Video and Dialogue Multimodal Fusion
-The multimodal fusion of text and video is very important when extracting factors, as it can provide multi-scale perspectives to more accurately extract factors from the teaching process.
-To better extract factors, we use Vision LLM to converse with children, or use conversations between teachers and children, along with video information during the process to extract key factors.
-The structure is shown in Figure 6.
-Figure 6: Important factors extraction based on video and conversation
-In this process, we can see that the Vision LLM-based multimodal algorithm can capture the corresponding states of children in various daily life scenarios and record these states in the form of dialogues.
-Ultimately, it obtains the corresponding factors through structured output.
-Specifically, in our software system, we use cameras to capture children’s facial expressions and engage in voice conversations with them, asking their opinions about the current task.
-Through multiple rounds of dialogue, we gather the essential features of children’s perceptions of specific subjects.
-Vision Large Language Model (V-LLM) is fully capable of performing this task.
-Because the V-LLM can communicate with children like a human and always remains patient.
-Additionally, the V-LLM provides emotional value feedback to children through storytelling and other methods, thereby attracting children to engage more willingly in communication.
-It is worth noting that the children’s voices will be converted into text and input into Vision LLM, while the responses generated by the Vision LLM will be converted into audio and played for the children to hear.
-Compared to simply recognizing facial expressions, this approach can provide individuals with a richer set of optional factors, thereby ensuring high-quality importance results.
-4.1.4 Physiological Data Analysis
-Electroencephalogram (EEG) data has been widely studied in the field of physiological data analysis.
-Since 2018, there has been a significant increase in the number of papers utilizing deep learning methods for emotion computation using EEG data.
-These methods include Convolutional Neural Networks (CNN), Deep Belief Networks (DBN), Recurrent Neural Networks (RNN), and Stacked Auto Encoders (SAE), among others.
-We will primarily handle data in modalities such as images and speech initially, and this type of data may be further studied in the future.
-4.1.5 A Multimodal Coding System for Student Behavior
-Multimodal sentiment analysis can lead to more accurate results.
-Common multimodal fusion techniques include feature-based fusion, model-based fusion, and decision-based late fusion.
-In future research, we will explore these three aspects according to the requirements of specific tasks.
-The current form of the multimodal encoding system for smart education is shown in Table LABEL:tab:coding.
-Table 1: Multimodal data coding system.
-Object Enviroment Multimodal data Learning Behavior Coding
-Qualitative
-characterization
-Student Intelligent recording system Sound data Audio Data Valid Active response S1 Proactive
-Passive response S2 Participation
-Ask questions
-proactively
-S3 Proactive
-Discussion S4 Proactive
-Sharing of results S5 Proactive
-Evaluation feedback S6 Proactive
-Invalid
-Ineffective confusion
-(topic deviation,
-intentional confusion)
-S7 Participation
-Silent Data Valid Think S8 Proactive
-Listening S9 Proactive
-Invalid Ineffective Silence S10 Participation
-Intelligent recording and broadcasting system image data Location Move Valid Seat M1 Participation
-Podium M2 Participation
-Between groups M3 Participation
-In group M4 Participation
-Invalid Corridor M5 Participation
-No location
-movement
-No location
-movement
-M6 Participation
-Body Movements Valid
-Look at the
-interactive whiteboard
-A1 Focus
-Watch the teacher A2 Focus
-Watch classmates A3 Focus
-Watch the
-learning tools
-A4 Focus
-Looking at
-the book
-A5 Focus
-For interactive
-whiteboard
-A6 Participation
-For teachers A7 Participation
-For companions A8 Participation
-For learning tools A9 Participation
-Book-oriented A10 Participation
-Invalid Not related to learning A11 Participation
-No body movements Sleep A12 Participation
-Distracted A13 Participation
-Learning platform (Wechat,MOOC etc.) text data Text data used by technology Valid technical support Sign in T1 Participation
-Answer T2 Participation
-Homework Exercises T3 Participation
-Self-learning T4 Participation
-Discussion T5 Participation
-Sharing of results T6 Participation
-Evaluation feedback T7 Participation
-Data collection T8 Participation
-Invalid technical support
-Unrelated to
-learning technology use
-T9 Participation
-No technology usage T10 Participation
-This will serve as a reference in the early stages of project development.
-4.2 Storing Various Factors for Different Children in a Graph Database
-To effectively organize the data, we extract key factors for each child individually and store these key factors in a graph database, as shown in Figure 7.
-Figure 7: Diagram of Storing Data Using a Graph Database
-In this data organization structure, we decouple time, attributes, and different children, making it more suitable for large-scale computation and processing.
-Additionally, this approach provides more opportunities to explore how to use knowledge graph-related technologies to investigate the deep relationships between habit formation in children and various factors.
 5 Identifying Important Factors in Education Process
@@
 Initially inspired, we found that we can use algorithms that identify important factors relevant to the field of habit formation to pinpoint factors that significantly impact education.
-We first introduce a classic method for calculating key factors based on the Growth Curve Model (Lewis et al.
-(2021)) and Shapley value (LEWIS et al.
-(2021)).
-Then, we propose a new algorithm based on Knowledge Graph methodologies to address the issues faced by these two methods.
-5.1 Growth Curve Model in Important Factors Identification
-Assuming we have 5 participants who record their habit automaticity scores (SRHI Automaticity) every day for 21 days.
-Additionally, we collect each participant’s initial commitment, previous mindfulness experience, and personality traits, including agreeableness and emotional stability.
-The collected sample format is illustrated in Table 2.
-Table 2: Table caption. For brevity, only a subset of the data is shown here
-Participant Day SRHI Automaticity Commitment Experience Agreeableness Emotional Stability
-1 1 2.0 High Low High Low
-1 2 2.2 High Low High Low
-1 21 3.5 High Low High Low
-2 1 1.5 Medium Medium Low High
-2 2 1.7 Medium Medium Low High
-5 21 4.2 Low High Medium Medium
-First, we aggregate the daily data of all participants into a long-format data frame, where each row represents a participant’s record on a specific day.
-Then we use a Growth Curve Model to analyze the data. The hierarchical structure of the model is as follows:
-  • Level-1 (Time Variation): Describes the change in individual habit automaticity over time.
-  • Level-2 (Individual Differences): Describes the differences between individuals, using time-invariant covariates such as initial commitment, experience, agreeableness, and emotional stability.
-The model equations can be represented as follows:
-$y_{ti}^{SRHI}A=b_{1i}+b_{2i}\cdot t+u_{ti}$ (1)
-where $b_{1i}=\beta_{01}+\beta_{11}\cdot Commitment_{i}+\beta_{21}\cdot Experience_{i}+\beta_{31}\cdot Agreeableness_{i}+\beta_{41}\cdot EmotionalStability_{i}+d_{1i}$, $b_{2i}=\beta_{02}+\beta_{12}\cdot Commitment_{i}+\beta_{22}\cdot Experience_{i}+\beta_{32}\cdot Agreeableness_{i}+\beta_{42}\cdot EmotionalStability_{i}+d_{2i}$.
-After obtaining the model equations, we can use Python to fit the above model and compute the corresponding parameters. The code is as follows:
-1import pandas as pd
-2import numpy as np
-3import statsmodels.api as sm
-4from statsmodels.formula.api import mixedlm
-5
-6np.random.seed(42)
-7days = np.arange(1, 22)
-8data = []
-9
-10for participant in range(1, 6):
-11 commitment = np.random.choice([’High’, ’Medium’, ’Low’])
-12 experience = np.random.choice([’High’, ’Medium’, ’Low’])
-13 agreeableness = np.random.choice([’High’, ’Medium’, ’Low’])
-14 emotional_stability = np.random.choice([’High’, ’Medium’, ’Low’])
-15
-16 for day in days:
-17 # Generate random SRHI Automaticity data (sample data)
-18 srhi_automaticity = np.random.normal(loc=3 + 0.1 * day, scale=0.5)
-19 data.append([participant, day, srhi_automaticity, commitment, experience, agreeableness, emotional_stability])
-20
-21df = pd.DataFrame(data, columns=[’participant’, ’day’, ’srhi_automaticity’, ’commitment’, ’experience’, ’agreeableness’, ’emotional_stability’])
-22
-23# Convert categorical variables to dummy variables
-24df = pd.get_dummies(df, columns=[’commitment’, ’experience’, ’agreeableness’, ’emotional_stability’], drop_first=True)
-25
-26# Define formula
-27formula = ’srhi_automaticity ~ day + commitment_Medium + commitment_Low + experience_Medium + experience_Low + agreeableness_Medium + agreeableness_Low + emotional_stability_Medium + emotional_stability_Low’
-28
-29# Fit a mixed-effects linear model
-30model = mixedlm(formula, df, groups=df[’participant’], re_formula=’~day’)
-31result = model.fit()
-32
-33# Output the model results summary
-34print(result.summary())
-After fitting the model, you can review the model parameters and significance levels. The fitted results are typically in the following form:
-1Mixed Linear Model Regression Results
-2=============================================================
-3Model: MixedLM Dependent Variable: srhi_automaticity
-4No. Observations: 105 Method: REML
-5No. Groups: 5 Scale: 0.2765
-6Min. group size: 21 Log-Likelihood: -95.2805
-7Max. group size: 21 Converged: Yes
-8Mean group size: 21.0
-9-------------------------------------------------------------
-10 Coef. Std.Err. z P>|z| [0.025 0.975]
-11-------------------------------------------------------------
-12Intercept 2.710 0.130 20.846 0.000 2.455 2.965
-13day 0.010 0.005 2.102 0.036 0.001 0.020
-14commitment_Medium -0.030 0.100 -0.300 0.764 -0.226 0.166
-15commitment_Low 0.100 0.120 0.833 0.405 -0.135 0.335
-16experience_Medium 0.030 0.010 3.000 0.003 0.010 0.050
-17experience_Low -0.030 0.020 -1.500 0.134 -0.070 0.010
-18agreeableness_Medium -0.020 0.010 -2.000 0.045 -0.040 0.000
-19agreeableness_Low 0.020 0.010 1.800 0.072 -0.002 0.042
-The new question is, how can we determine which factors have a significant impact on habit formation from this result?
-The answer lies in observing the regression coefficients and significance levels.
-Specifically, in the model results summary, we focus on each covariate’s regression coefficient (Coef.), standard error (Std.
-Err.), z-value (z), and p-value (P $>|z|$).
-The p-value is used to determine the significance of each covariate’s impact on the change in habit automaticity.
-Typically, we use the following criteria to judge significance:
-  • p $<$ 0.05: The covariate has a significant effect on the outcome variable.
-  • p $<$ 0.01: The covariate has a highly significant effect on the outcome variable.
-  • p $<$ 0.1: The covariate has a marginally significant effect on the outcome variable.
-For significant covariates, the regression coefficient indicates the direction and magnitude of that variable’s effect on habit automaticity scores. For example:
-  • A positive regression coefficient indicates that an increase in the covariate leads to an increase in habit automaticity scores.
-  • A negative regression coefficient indicates that an increase in the covariate leads to a decrease in habit automaticity scores.
-Therefore, we can analyze the model output as follows:
-  • Time Effect (day):
-    Regression coefficient is 0.01, p-value is 0.036 (p ¡ 0.05), indicating that habit automaticity significantly increases over time.
-  • Commitment:
-    commitment Medium: Regression coefficient is -0.03, p-value is 0.764 (p ¿ 0.05), indicating that medium commitment has no significant effect on the change in habit automaticity.
-    commitment Low: Regression coefficient is 0.10, p-value is 0.405 (p ¿ 0.05), indicating that low commitment has no significant effect on the change in habit automaticity.
-  • Mindfulness Experience (Experience):
-    experience Medium : Regression coefficient is 0.03, p-value is 0.003 (p ¡ 0.01), indicating that medium mindfulness experience has a highly significant positive effect on the increase in habit automaticity.
-    experience Low: Regression coefficient is -0.03, p-value is 0.134 (p ¿ 0.05), indicating that low mindfulness experience has no significant effect on habit automaticity.
-  • Agreeableness:
-    ‘agreeableness Medium: Regression coefficient is -0.02, p-value is 0.045 (p ¡ 0.05), indicating that medium agreeableness has a significant negative effect on habit automaticity.
-    agreeableness Low: Regression coefficient is 0.02, p-value is 0.072 (p ¡ 0.1), indicating that low agreeableness has a marginally significant positive effect on habit automaticity.
-  • Emotional Stability:
-    emotional stability Medium‘: Regression coefficient is 0.02, p-value is 0.072 (p ¡ 0.1), indicating that medium emotional stability has a marginally significant positive effect on habit automaticity.
-    emotional stability Low: Regression coefficient is 0.02, p-value is 0.072 (p ¡ 0.1), indicating that low emotional stability has a marginally significant positive effect on habit automaticity.
-Thus, we can conclude that the factors significantly affecting habit formation are as follows:
-  • Time Effect (day): Habit automaticity significantly increases over time.
-  • Mindfulness Experience (Experience): Medium mindfulness experience significantly promotes an increase in habit automaticity.
-  • Agreeableness: Medium agreeableness significantly inhibits the increase in habit automaticity.
-  • Emotional Stability: Both medium and low emotional stability have a marginally significant positive effect on habit automaticity.
-Through this analytical approach, we can identify which individual characteristics significantly influence the speed of habit formation, thereby providing a basis for personalized interventions.
-5.2 Shapely Algorithm in Important Factors Identification
-LEWIS et al. (2021) using the Shapley algorithm (Winter (2002)), important factors influencing habit formation are extracted from survey data. This process mainly includes the following steps:
-  • Data Collection: Collect data from participants’ daily surveys, including whether they practiced breathing exercises, the level of automation, sense of reward, sense of challenge, confidence, motivation, mood, and daily activities.
-  • Model Training: Use the collected data to train a binary classification model (e.g., Random Forest) to predict whether participants will practice breathing exercises the next day.
-  • Shapley Value Calculation: For each model prediction, calculate the Shapley value for each feature, measuring its contribution to the prediction result.
-Shapley values represent the marginal contribution of features to the prediction result.
-  • Feature Importance Analysis: By aggregating Shapley values, determine which features are most important for model predictions.
-In summary, the Shapley method is more direct compared to the Growth Curve Model.
-The Growth Curve Model requires calculating significant statistical information from long-term surveys to determine the important factors influencing habit formation.
-In contrast, Shapley can calculate this importance by processing a single day’s data, and further improves prediction accuracy by accumulating data over multiple days.
-Below is a simplified Python implementation process, demonstrating how to use the Shapley value algorithm to extract important features from constructed data.
-First, we need to construct a simulated dataset. Let’s assume we have the following features:
-  • did breathing: Whether the breathing exercise was performed (0 or 1)
-  • automaticity: Level of automation (rating from 1 to 7)
-  • rewarding: Sense of reward (rating from 1 to 7)
-  • challenging: Sense of challenge (rating from 1 to 7)
-  • confidence: Confidence (rating from 1 to 7)
-  • motivation: Motivation (rating from 1 to 7)
-  • mood: Mood (rating from 1 to 7)
-  • busy: Level of busyness (rating from 1 to 7)
-These data can be implemented and processed using the following Python code:
-1import numpy as np
-2import pandas as pd
-3from sklearn.ensemble import RandomForestClassifier
-4from sklearn.model_selection import train_test_split
-5import shap
-6
-7np.random.seed(42)
-8data_size = 1000
-9data = pd.DataFrame({
-10 ’did_breathing’: np.random.randint(0, 2, data_size),
-11 ’automaticity’: np.random.randint(1, 8, data_size),
-12 ’rewarding’: np.random.randint(1, 8, data_size),
-13 ’challenging’: np.random.randint(1, 8, data_size),
-14 ’confidence’: np.random.randint(1, 8, data_size),
-15 ’motivation’: np.random.randint(1, 8, data_size),
-16 ’mood’: np.random.randint(1, 8, data_size),
-17 ’busy’: np.random.randint(1, 8, data_size)
-18})
-19
-20# Features and Labels
-21X = data.drop(’did_breathing’, axis=1)
-22y = data[’did_breathing’]
-23
-24# Traning and Validation datasets
-25X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2, random_state=42)
-26
-27# Train Random Forest Classifier
-28model = RandomForestClassifier(n_estimators=100, random_state=42)
-29model.fit(X_train, y_train)
-30
-31# Using SHAP to Explain the Model
-32explainer = shap.TreeExplainer(model)
-33shap_values = explainer.shap_values(X_test)
-34
-35# Features Importance Visualization
-36shap.summary_plot(shap_values[1], X_test)
-37
-38# Print Absolute Averaged Shapley Value
-39shap_importance = np.abs(shap_values[1]).mean(axis=0)
-40feature_importance = pd.DataFrame(
-41 list(zip(X_train.columns, shap_importance)),
-42 columns=[’Feature’, ’Importance’]
-43).sort_values(by=’Importance’, ascending=False)
-44
-45print(feature_importance)
-Through the above process, we can determine which features are most important to the model’s predictions, thereby understanding which factors significantly influence habit formation.
-5.3 Knowledge Graph Algorithms in Important Factors Identification
-The Growth Curve Model and Shapley value method both employ very simple linear fitting or simple classification models, which are very effective when dealing with small amounts of data (e.g., 62 participants).
-However, such algorithms cannot handle large-scale data.
-Additionally, these algorithms do not consider the differences between individuals; important factors are directly derived from the averages of statistical data, assuming these factors are important for everyone.
-This is intuitively unreasonable.
-In contrast, Graph Neural Networks (GNNs) can be used to handle complex relationships and structured data, enabling them to capture relationships within the data better than traditional machine learning models in some cases.
-To better distinguish individual differences and conduct habit formation research on large-scale data, we first store all children and related factors in a graph database (we chose Neo4J, which can be easily interacted with using Python).
-Specific details can be found in Chapter 2.3.
-Inspired by the work of LEWIS et al.
-(2021), we use Graph Neural Networks (such as Graph Convolutional Networks, GCN (Zhang et al.
-(2019)), or Graph Attention Networks, GAT (Veličković et al.
-(2017))) to train the model.
-The model can propagate information through the graph via a message-passing mechanism, capturing more complex relationships.
-The goal of the GNN model is to predict whether each participant will perform a breathing exercise next time.
-Although GNNs are inherently black-box models, techniques similar to Shapley values (such as GNNExplainer Ying et al.
-(2019)) can be used to explain the contribution of each node and edge to the final prediction.
-GNNExplainer optimizes a subgraph to maximize the explanation of the model’s prediction, thereby identifying which nodes and edges are most important for the prediction.
 6 Regulating Children’s Behavior Using Important Factors
@@
 One classic experiment is the reward devaluation experiment, where animals reduce behaviors related to rewards that are no longer attractive (for example, when food is poisoned).
-The habit system refers to the behavior system triggered automatically by specific situations or stimuli, rather than controlled by the consequences of behavior.
-This system is formed through repetition and reinforcement, making behavior automatic and not requiring consideration of consequences.
-Habit behavior primarily involves the brain’s dorsolateral striatum (DLS).
-These areas are related to motor control and automated behaviors.
-The habit system is more like a stress response, such as stopping at a red light every time you see it.
-The goal-directed system and the habit system represent two extreme forms of behavior control.
-The goal-directed system allows individuals to flexibly adjust their behavior to achieve optimal outcomes, while the habit system automates behavior, thereby conserving cognitive resources.
-In real life, most behaviors result from the combined action of these two systems, dynamically adjusting between them based on context and experience.
-In recent years, reinforcement learning has gradually gained attention and has been introduced into the study of behavior control (Fan et al.
-(2023); Wu and Wang (2023)).
-In the following content, I will discuss in detail the current theoretical models of behavior control and, based on these models, propose my own framework for regulating children’s behavior by incorporating important factors.
-6.1 Binding and Retrieval in Action Control (BRAC)
-6.1.1 Basic Concepts
-Frings et al.
-(2020) believe that human behavior control relies on representations that integrate perception and action.
-By associating specific behaviors with their outcomes, humans learn to act purposefully.
-Based on this cognitive hypothesis, Frings et al.
-(2020) proposed the BRAC framework, aimed at integrating and explaining a wide array of research findings on human behavior control.
-Specifically, the BRAC framework proposes two core processes: feature binding and retrieval.
-Binding refers to integrating the features of stimuli, responses, and effects into an event file; retrieval refers to automatically retrieving previously bound event files when a particular feature is repeated.
-The framework emphasizes that binding and retrieval are functionally distinct and separable processes, and these two processes are independently regulated by top-down and bottom-up processes.
-The BRAC framework can integrate and explain phenomena of behavior control across various experimental paradigms, such as task switching, negative priming, and repetition priming.
-Through this framework, different behavior control tasks can be reinterpreted as outcomes driven by binding and retrieval processes.
-In addition, BRAC can also be extended to other related fields such as attention, memory and learning, and motivation.
-For example, in attention research, feature binding and retrieval can explain cross-trial effects in spatial orientation and visual search tasks; in memory and learning research, the concept of event files is closely related to episodic memory; in motivation research, BRAC can explain how approach and avoidance behaviors influence the evaluation of neutral stimuli.
-The BRAC framework helps to unify and simplify theoretical explanations of behavior control phenomena, facilitating the integration of results across different experimental paradigms.
-The framework also raises new hypotheses and research questions, such as how to distinguish the specific contributions of binding and retrieval processes in behavior control, and what the neural bases of these processes are.
-6.1.2 Children’s Behavior Regularization based on Important Factors and BRAC
-The BRAC theory emphasizes two core processes in behavior control: feature binding and retrieval.
-Inspired by this theory, we can utilize it to influence behavior by changing the processes of feature binding and retrieval of important factors.
-Specifically, we can design behavior interventions in three aspects: enhancing positive feature binding, reducing negative feature binding, and increasing the effectiveness of feature retrieval.
-In enhancing positive feature binding, we can strengthen the positive feature binding associated with target behavior through positive experiences and feedback.
-For example, we can design reward mechanisms that offer incentives for completing habit practices, such as points, badges, or small gifts, to enhance the sense of reward from practice.
-Additionally, we can use LLMs to provide instant positive feedback, such as encouraging words or progress charts, to increase the enjoyment and sense of accomplishment from practice.
-In reducing negative feature binding, we can minimize negative experiences and emotional bindings associated with target habits.
-For instance, we can lower the difficulty of habit practice by breaking tasks down or providing guidance, making them easier to complete.
-We can also introduce relaxation techniques or time management strategies to help individuals better cope with stress and busy schedules.
-In increasing the effectiveness of feature retrieval, we can enhance the retrieval process of relevant features through environmental cues and behavior prompts, promoting the automation of habits.
-For example, we can set reminders at key times and locations, such as phone notifications or calendar alerts, to remind individuals to practice their habits.
-We can also bind target habits to existing daily behaviors, such as meditating after brushing teeth, leveraging the automation of existing habits to trigger new ones.
-In addition to these interventions, we can design personalized interventions based on the characteristics and needs of different individuals.
-For example, adjusting the intensity and content of interventions based on different emotional states and levels of busyness.
-Specifically, during low emotional states, we can try increasing positive feedback and rewards, providing more emotional support.
-During busy states, we can simplify the steps of habit practice and offer time management suggestions.
-By continuously collecting individual feedback and behavior data, we can dynamically adjust interventions to ensure their continued effectiveness.
-For instance, using smartphone applications to track individuals’ habit practice status and adjusting reminders and feedback strategies based on the data.
-6.2 Optimal Behavior Estimation by Reforcement Learning
-BRAC has conducted systematic research on the role of stress responses in individual behavior, but it has yet to address an important issue: what specific actions should an individual take when they are in a certain state?
-For instance, when a person is sad, they might need to talk to someone or silently endure.
-This decision depends on the individual’s factors, such as personality.
-However, how to use a behavior intervention system to help individuals overcome negative emotions and make optimal behavioral decisions is still a question worth exploring.
-Furthermore, how to utilize reinforcement learning to determine the optimal behavior, thereby achieving positive goal (such as efficient habit formation) is an area worthy of research.
-To this end, a feasible solution is to first use reinforcement learning to build a behavior model.
-The output of this model would be the action the user should take in their current state, and this action would always tend to obtain more rewards (e.g., it would more easily help form good habits).
-The overall framework of this system is shown in Figure 8.
-Figure 8: Motivation for Optimal Behavior Estimation Based on RL
-In this system, we will use the techniques from Chapters 2 and 3 to extract important factors as state inputs for the reinforcement learning model.
-For a preference-trained reinforcement learning model, the output will be more inclined to produce actions that align with the user’s preferences (favoring the pursuit of a better emotional state).
-6.2.1 Behavioral Network based on Reforcement Learning
-Yamada and Toda (2023) proposed a Q-learning method that can effectively help us understand how to use reinforcement learning to model the learning of behaviors and the corresponding decision-making processes.
-Reinforcement learning is used to simulate the habit formation process within a behavior network.
-This study assumes that behavior is a network composed of multiple responses, with reinforcement learning used to represent the agent’s history of rewards obtained during transitions between different responses.
-The agent uses the Q-learning algorithm to learn the Q-values (i.e., action values) between various responses, thereby guiding behavior selection and path searching.
-The agent updates the Q-values using the Q-learning algorithm, where $Q(a_{t-1},a_{t})$ represents the action value of transitioning from response $a_{t-1}$ to response $a_{t}$.
-The Q-value update formula is:
-$Q(a_{t-1},a_{t})\leftarrow Q(a_{t-1},a_{t})+\alpha\delta$ (2)
-where $\alpha$ is the learning rate and $\delta$ is the reward prediction error. The formula for calculating the reward prediction error $\delta$ is:
-$\delta=R(a_{t})+\gamma\max_{a_{t+1}}Q(a_{t},a_{t+1})-Q(a_{t-1},a_{t})$ (3)
-where $R(a_{t})$ is the reward obtained when transitioning to response $a_{t}$, and $\gamma$ is the discount factor for future rewards.
-The agent selects responses based on Q-values through a softmax function, which increases the probability of selecting responses with higher Q-values.
-The agent also uses shortest path search algorithms (such as Dijkstra’s algorithm) to find the shortest path from the current response to the target response in the behavior network, thereby optimizing the behavior sequence.
-By simulating different training amounts, reward schedules, and selection environments, the study demonstrated how the reinforcement learning model can reproduce the habit formation phenomena observed in experiments.
-The simulation results indicate that habit formation can be viewed as a change in the structure of the behavior network rather than behavior controlled by two separate systems.
-This work helps researchers establish a unified theoretical framework to explain the transition process between goal-directed behavior and habitual behavior.
-6.2.2 Preference-Aligned Reinforcement Learning Algorithm
-Inspired by the work of Yamada and Toda (2023) and others, we propose a preference-aligned reinforcement learning algorithm. As illustrated in Figure 9.
-Figure 9: Preference-Aligned Reinforcement Learning Algorithm
-First, a standard supervised learning method is used to train an initial model.
-Typically, this initial model is trained on a large-scale dataset to ensure it has basic task capabilities.
-Then, in our designed multimodal human-computer interaction system, the model is further refined using human feedback to make its outputs more inclined to produce positive results.
-Specifically, the model generates multiple candidate outputs for a given input.
-In our interactive system, human evaluators rank or score these candidate outputs, and the optimal output is selected based on preferences that lead to positive outcomes.
-A reward model is trained using the human feedback data. This reward model learns how to evaluate the quality of the model outputs based on human preferences.
-Furthermore, the trained reward model is used as a reward signal to further optimize the initial model through reinforcement learning algorithms (e.g., policy gradient methods).
-Common steps include using feedback from the reward model to update the policy, making the model more inclined to generate high-scoring outputs, and continuously evaluating the policy’s performance and optimizing it based on the feedback.
-This process can be iterated, collecting feedback, training the reward model, and updating the policy in multiple rounds to continuously enhance the model’s performance and alignment.
-7 Conclusion
-In our implementation plan, the multimodal interaction-based system is divided into the following three parts:
-  • Dynamic factor extraction based on artificial intelligence algorithms
-  • Identification of key factors
-  • Child behavior regulation system
-We intend to build a comprehensive and closed-loop preschool education interaction system from these three aspects.
-The required technologies and their respective completion statuses are shown in the table below.
-Table 3: The technical points required to implement the current system and their progress
-Required Item Application
-Current
-Status
-Factor Discovery System
-Smart
-Home
-Hardware
-Collect multimodal interaction information
-of children in a fixed room
-0%
-Software Interaction Child identification
-To identify the identity of the child currently
-interacting
-80%
-Emotion Analysis Extracting emotional factors 80%
-Digital Human
-Communicate with the child, provide emotional
-value, and extract structured factor information
-90%
-Entertainment System Improve user stickiness 0%
-HCI Plugins
-Extracting children’s behavioral factors in
-third-party software
-0%
-Graph Storage
-Using graph databases to store
-structured information
-100%
-Important Factor Identification Growth Curve Data Collection Enabled by Factor Discovery System 50%
-Analysis Method
-Methods for analyzing data and
-identifying important factors
-100%
-Shapely Data Collection Enabled by Factor Discovery System 50%
-Analysis Method
-Methods for analyzing data and
-identifying important factors
-100%
-Knowledge Graph Data Collection Enabled by Factor Discovery System 50%
-Analysis Method
-Methods for analyzing data and
-identifying important factors
-60%
-Child Behavior Regulation System
-Behavior Regularization
-based on
-Important Factors and BRAC
-Regulating children’s behavior 0%
-Optimal Behavior Estimation by RL
-RL in
-behavior modeling
-Behavioral Model 80%
-PARL
-Preference-aligned reinforcement
-learning algorithm
-10%
-Based on the current progress, we predict that such a system can be implemented within two years, and the required personnel will not exceed two people (including myself).
-The contributions of the proposed solution are as follows.
-  • Proposed a complete multimodal interactive education system.
-  • Divided the application of AI in the educational process into three parts and accordingly proposed innovative algorithms and solutions.